OpenStudy (anonymous):

find the x-intercepts of y=-x2-4x+12

OpenStudy (anonymous):

you need to factor the trinomial - can you do this?

OpenStudy (anonymous):

oooh - it wont factor so you need to use the quadratic formula

OpenStudy (anonymous):

firsly you ley y=0, so you have 0=-x^2-4x+12, multiply by -1 through out 0=x^2+4x-12 0=(x+6)(x-2) x=-6 or x=2, therefore your x intercepts are (-6;0) and (2;0)

OpenStudy (anonymous):

right sihle - i misread the question - didn't notice the - x!!

OpenStudy (anonymous):

-x^2